OpenStudy (anonymous):

let f(x) be a function satisfying f(x)f(y)=f(xy) for all real x,y . If f(2) = 4 ,then what is the value of f(1/2)?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

substitute x=y=1 : f(1)*f(1)-f(1)=0 f(1)=0 or f(1) =1 ----------> [1] substitute x=2 : f(2)*f(1/2)=f(1); f(1/2)=f(1) / 4 = either 0 or 1/4

OpenStudy (anonymous):

but f(2)f(1)=f(2) f(1)=1 --- > [2] hence, the only answer is f(1/2) = 1/4